A gas station is providing the state auto inspection service for the general public. As soon as the gas station opens at 7AM, cars arrive for inspection following a Poisson process with rate λ (arrivals/minute). (a) Assume that each inspection takes a constant amount time, namely c (minutes). Let W2 denote the random time that the second vehicle waits between arriving and the time at which its inspection starts. What is the probability that W2 = 0? What is the expected value of W2? (b) Answer the same questions above assuming that the service time for the first customer is an exponentially distributed random variable with mean 1/k
Answer #1

Let x be the interarrival time between ith arriving car and (i - 1)th and { x } be the independent variable.

if x2\geq c, the second arriving car need not do wait, then we have

P { x2\geq c } e-\lambdac

if x2\propto c, it needs to wait (c - x2) minutes,

Let x2 = t, then we have

E (c - t) = \int_{0}^{c}(c - t) \lambda e-\lambdat dt + \int_{c}^{+\infty } \lambda e-\lambdat dt

Let y be the inspection time for the first arriving car.

using a condition, Y = y

if x2\geq y, the second will not need to wait. otherwise, it will wait (y - t) minutes.

So, the probability that it will not wait is given by :

P { x2\geq Y } = \int_{c}^{+\infty } P { x2\geq Y | Y

P { x2\geq Y } = y \int_{c}^{+\infty } f (y) dy

P { x2\geq Y } = \int_{c}^{+\infty } e-\lambday\mu e-\muy dy

P { x2\geq Y } = \mu / (\lambda + \mu)

And the mean waiting time which will be given as :

P { x2\geq Y } = \int_{c}^{+\infty } \int_{0}^{y}(y - t) \lambda e-\lambdat dt . \mu e-\muy dy

P { x2\geq Y } = \int_{c}^{+\infty } [y + (1 / \lambda) e-\lambday - (1 / \lambda)] \mu e-\muy dy

P { x2\geq Y } = \lambda / \mu (\lambda + \mu)